OpenStudy (anonymous):

A fixed system of charges exerts a force of magnitude 21 N on a 6 C charge. The 6 C charge is replaced with a 8 C charge. What is the exact magnitude of the force (in N) exerted by the system of charges on the 8 C charge?

OpenStudy (anonymous):

f=(k*q1*q2)/r^2

OpenStudy (anonymous):

you keep everything except one of the q's constant i would do this 21=6kq/r^2 21/6=kq/r^2 then multiply by 8 and i got 28

OpenStudy (anonymous):

does that make sense? i was kinda messy

OpenStudy (anonymous):

I think. What did you use for your r?